OpenStudy (anonymous):

a lil trouble with lcd's Find LCD : 1/8 + 1/4 + 1/3 + 1/6

OpenStudy (anonymous):

dont give me the answer

OpenStudy (anonymous):

help me

OpenStudy (anonymous):

what would all of those numbers have in common

Parth (parthkohli):

Find the LCM of 8,4,3,6 in other words.

OpenStudy (amistre64):

i tend to just start clearing fractions by multiplying thru by the smallest denominator

OpenStudy (jiteshmeghwal9):

find the L.C.M of all the numbers in the denominator

Parth (parthkohli):

\( \color{Black}{\Rightarrow 8 = 2 \times 2 \times 2}\) \( \color{Black}{\Rightarrow 4 = 2 \times 2}\) \( \color{Black}{\Rightarrow 3 = 3}\) \( \color{Black}{\Rightarrow 6 = 3 \times 2}\) \(L.C.M = 2 \times 2 \times 2 \times3\)

OpenStudy (jiteshmeghwal9):

or u can u can solve them through grouping